# YOLOv10: Real-Time End-to-End Object Detection

### v1

@@ Abstract
Over the past years, YOLOs have emerged as the predominant paradigm in the field of real-time object detection owing to their effective balance between computational cost and detection performance. Researchers have explored the architectural designs, optimization objectives, data augmentation strategies, and others for YOLOs, achieving notable progress. However, the reliance on the non-maximum suppression (NMS) for post-processing hampers the end-to-end deployment of YOLOs and adversely impacts the inference latency. Besides, the design of various components in YOLOs lacks the comprehensive and thorough inspection, resulting in noticeable computational redundancy and limiting the model’s capability. It renders the suboptimal efficiency, along with considerable potential for performance improvements. In this work, we aim to further advance the performance-efficiency boundary of YOLOs from both the post-processing and the model architecture. To this end, we first present the consistent dual assignments for NMS-free training of YOLOs, which brings the competitive performance and low inference latency simultaneously. Moreover, we introduce the holistic efficiency-accuracy driven model design strategy for YOLOs. We comprehensively optimize various components of YOLOs from both the efficiency and accuracy perspectives, which greatly reduces the computational overhead and enhances the capability. The outcome of our effort is a new generation of YOLO series for real-time end-to-end object detection, dubbed YOLOv10. Extensive experiments show that YOLOv10 achieves the state-of-the-art performance and efficiency across various model scales. For example, our YOLOv10-S is 1.8 faster than RT-DETR-R18 under the similar AP on COCO, meanwhile enjoying 2.8 smaller number of parameters and FLOPs. Compared with YOLOv9-C, YOLOv10-B has 46% less latency and 25% fewer parameters for the same performance. Code: https://github.com/THU-MIG/yolov10.
@@ 1 Introduction
Specifically, YOLOs usually employ one-to-many label assignment strategy during training, whereby one ground-truth object corresponds to multiple positive samples. Despite yielding superior performance, this approach necessitates NMS to select the best positive prediction during inference. This slows down the inference speed and renders the performance sensitive to the hyperparameters of NMS, thereby preventing YOLOs from achieving optimal end-to-end deployment [71]. One line to tackle this issue is to adopt the recently introduced end-to-end DETR architectures [4, 74, 67, 28, 34, 40, 61]. For example, RT-DETR [71] presents an efficient hybrid encoder and uncertainty-minimal query selection, propelling DETRs into the realm of real-time applications. Nevertheless, the inherent complexity in deploying DETRs impedes its ability to attain the optimal balance between accuracy and speed. Another line is to explore end-to-end detection for CNN-based detectors, which typically leverages one-to-one assignment strategies to suppress the redundant predictions [5, 49, 60, 73, 16]. However, they usually introduce additional inference overhead or achieve suboptimal performance.
@@ 2 Related Work
End-to-end object detectors. End-to-end object detection has emerged as a paradigm shift from traditional pipelines, offering streamlined architectures [48]. DETR [4] introduces the transformer architecture and adopts Hungarian loss to achieve one-to-one matching prediction, thereby eliminating hand-crafted components and post-processing. Since then, various DETR variants have been proposed to enhance its performance and efficiency [40, 61, 50, 28, 34]. Deformable-DETR [74] leverages multi-scale deformable attention module to accelerate the convergence speed. DINO [67] integrates contrastive denoising, mix query selection, and look forward twice scheme into DETRs. RT-DETR [71] further designs the efficient hybrid encoder and proposes the uncertainty-minimal query selection to improve both the accuracy and latency. Another line to achieve end-to-end object detection is based CNN detectors. Learnable NMS [23] and relation networks [25] present another network to remove duplicated predictions for detectors. OneNet [49] and DeFCN [60] propose one-to-one matching strategies to enable end-to-end object detection with fully convolutional networks. FCOS [73] introduces a positive sample selector to choose the optimal sample for prediction.
@@ 4.3 Model Analyses
1 | base. | 11.2/25.9 | 28.6/78.9 | 44.3/50.3 | 2.44/5.22 |
2 | +cls. | 9.9/23.2 | 23.5/67.7 | 44.2/50.2 | 2.39/5.07 |
3 | +downs. | 8.0/19.7 | 22.2/65.0 | 44.4/50.4 | 2.36/4.97 |
4 | +block. | 6.2/14.1 | 20.8/58.1 | 44.5/50.4 | 2.31/4.57 |
@@ A.1 Implementation Details
Moreover, we follow [71] to establish the end-to-end speed benchmark. Since the execution time of NMS is affected by the input, we thus measure the latency on the COCO val set, like [71]. We adopt the same NMS hyperparameters used by the detectors during their validation. The TensorRT efficientNMSPlugin is appended for post-processing and the I/O overhead is omitted. We report the average latency across all images.
@@ A.3 Details of Rank-Guided Block Design
A.4 More Results on COCO
We report the detailed performance of YOLOv10 on COCO, including AP and AP at different IoU thresholds, as well as AP, AP, and AP across different scales, in Tab. 15.
A.5 More Analyses for Holistic Efficiency-Accuracy Driven Model Design
@@
A.6 Visualization Results
@@
A.7 Contribution, Limitation, and Broader Impact

### v2

@@ Abstract
Over the past years, YOLOs have emerged as the predominant paradigm in the field of real-time object detection owing to their effective balance between computational cost and detection performance. Researchers have explored the architectural designs, optimization objectives, data augmentation strategies, and others for YOLOs, achieving notable progress. However, the reliance on the non-maximum suppression (NMS) for post-processing hampers the end-to-end deployment of YOLOs and adversely impacts the inference latency. Besides, the design of various components in YOLOs lacks the comprehensive and thorough inspection, resulting in noticeable computational redundancy and limiting the model’s capability. It renders the suboptimal efficiency, along with considerable potential for performance improvements. In this work, we aim to further advance the performance-efficiency boundary of YOLOs from both the post-processing and the model architecture. To this end, we first present the consistent dual assignments for NMS-free training of YOLOs, which brings the competitive performance and low inference latency simultaneously. Moreover, we introduce the holistic efficiency-accuracy driven model design strategy for YOLOs. We comprehensively optimize various components of YOLOs from both the efficiency and accuracy perspectives, which greatly reduces the computational overhead and enhances the capability. The outcome of our effort is a new generation of YOLO series for real-time end-to-end object detection, dubbed YOLOv10. Extensive experiments show that YOLOv10 achieves the state-of-the-art performance and efficiency across various model scales. For example, our YOLOv10-S is 1.8 faster than RT-DETR-R18 under the similar AP on COCO, meanwhile enjoying 2.8 smaller number of parameters and FLOPs. Compared with YOLOv9-C, YOLOv10-B has 46% less latency and 25% fewer parameters for the same performance. Code and models are available at https://github.com/THU-MIG/yolov10.
@@ 1 Introduction
Specifically, YOLOs usually employ one-to-many label assignment strategy during training, whereby one ground-truth object corresponds to multiple positive samples. Despite yielding superior performance, this approach necessitates NMS to select the best positive prediction during inference. This slows down the inference speed and renders the performance sensitive to the hyperparameters of NMS, thereby preventing YOLOs from achieving optimal end-to-end deployment [78]. One line to tackle this issue is to adopt the recently introduced end-to-end DETR architectures [4, 81, 73, 30, 36, 42, 67]. For example, RT-DETR [78] presents an efficient hybrid encoder and uncertainty-minimal query selection, propelling DETRs into the realm of real-time applications. Nevertheless, when considering only the forward process of model during deployment, the efficiency of the DETRs still has room for improvements compared with YOLOs. Another line is to explore end-to-end detection for CNN-based detectors, which typically leverages one-to-one assignment strategies to suppress the redundant predictions [6, 55, 66, 80, 17]. However, they usually introduce additional inference overhead or achieve suboptimal performance for YOLOs.
@@ 2 Related Work
End-to-end object detectors. End-to-end object detection has emerged as a paradigm shift from traditional pipelines, offering streamlined architectures [53]. DETR [4] introduces the transformer architecture and adopts Hungarian loss to achieve one-to-one matching prediction, thereby eliminating hand-crafted components and post-processing. Since then, various DETR variants have been proposed to enhance its performance and efficiency [42, 67, 56, 30, 36, 28, 5, 77, 82]. Deformable-DETR [81] leverages multi-scale deformable attention module to accelerate the convergence speed. DINO [73] integrates contrastive denoising, mix query selection, and look forward twice scheme into DETRs. RT-DETR [78] further designs the efficient hybrid encoder and proposes the uncertainty-minimal query selection to improve both the accuracy and latency. Another line to achieve end-to-end object detection is based CNN detectors. Learnable NMS [24] and relation networks [26] present another network to remove duplicated predictions for detectors. OneNet [55] and DeFCN [66] propose one-to-one matching strategies to enable end-to-end object detection with fully convolutional networks. FCOS [80] introduces a positive sample selector to choose the optimal sample for prediction.
@@ 3.1 Consistent Dual Assignments for NMS-free Training
Discussion with other counter-parts. Similarly, previous works [28, 5, 77, 54, 6, 82, 45] explore the different assignments to accelerate the training convergence and improve the performance for different networks. For example, H-DETR [28], Group-DETR [5], and MS-DETR [77] introduce one-to-many matching in conjunction with the original one-to-one matching by hybrid or multiple group label assignments, to improve upon DETR. Differently, to achieve the one-to-many matching, they usually introduce extra queries or repeat ground truths for bipartite matching, or select top several queries from the matching scores, while we adopt the prediction aware assignment that incorporates the spatial prior. Besides, LRANet [54] employs the dense assignment and sparse assignment branches for training, which all belong to the one-to-many assignment, while we adopt the one-to-many and one-to-one branches. DEYO [45, 43, 44] investigates the step-by-step training with one-to-many matching in the first stage for convolutional encoder and one-to-one matching in the second stage for transformer decoder, while we avoid the transformer decoder for end-to-end inference. Compared with works [6, 80] which incorporate dual assignments for CNN-based detectors, we further analyze the supervision gap between the two heads and present the consistent matching metric for YOLOs to reduce the theoretical supervision gap. It improves performance through better supervision alignment and eliminates the need for hyper-parameter tuning.
@@ 4.3 Model Analyses
| 1 | base. | 11.2/25.9 | 28.6/78.9 | 44.3/50.3 | 2.44/5.22 |
| 2 | +cls. | 9.9/23.2 | 23.5/67.7 | 44.2/50.2 | 2.39/5.07 |
| 3 | +downs. | 8.0/19.7 | 22.2/65.0 | 44.4/50.4 | 2.36/4.97 |
| 4 | +block. | 6.2/14.1 | 20.8/58.1 | 44.5/50.4 | 2.31/4.57 |
@@
Performance gap compared with one-to-many training. Although achieving superior end-to-end performance under NMS-free training, we observe that there still exists the performance gap compared with the original one-to-many training using NMS, as shown in Tab. 5 and Tab. 1. Besides, we note that the gap diminishes as the model size increases. Therefore, we reasonably concludes that such a gap can be attributed to the limitations in the model capability. Notably, unlike the original one-to-many training using NMS, the NMS-free training necessitates more discriminative features for one-to-one matching. In the case of the YOLOv10-N model, its limited capacity results in extracted features that lack sufficient discriminability, leading to a more notable performance gap of 1.0% AP. In contrast, the YOLOv10-X model, which possesses stronger capability and more discriminative features, shows no performance gap between two training strategies. In Fig. 4, we visualize the average cosine similarity of each anchor point’s extracted features with those of all other anchor points on the COCO val set. We observe that as the model size increases, the feature similarity between anchor points exhibits a downward trend, which benefits the one-to-one matching. Based on this insight, we will explore approaches to further reduce the gap and achieve higher end-to-end performance in the future work.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by National Natural Science Foundation of China (Nos. 61925107, 62271281) and Beijing Natural Science Foundation (No. L223023).
@@ A.1 Implementation Details
Moreover, we follow [78] to establish the end-to-end speed benchmark. Since the execution time of NMS is affected by the input, we thus measure the latency on the COCO val set with the batch size of 1, like [78]. We adopt the same NMS hyperparameters used by the detectors during their validation. The TensorRT efficientNMSPlugin is appended for post-processing and the I/O overhead is omitted. We report the average latency across all images.
@@ A.3 Details of Rank-Guided Block Design
| Model | Epoch | Speed (epoch/hour) | Time (hour) |
| YOLOv6-3.0-M | 300 | 7.2 | 41.7 |
| YOLOv8-M | 500 | 18.3 | 27.3 |
| YOLOv9-M | 500 | 12.3 | 40.7 |
| Gold-YOLO-M | 300 | 4.7 | 63.8 |
| YOLO-MS | 300 | 7.1 | 42.3 |
| YOLOv10-M-o2o | 500 | 18.8 | 26.7 |
| YOLOv10-M | 500 | 17.2 | 29.1 |
| Model | Latency |
| YOLOv10-N | 6.19ms |
| YOLOv10-S | 7.15ms |
| YOLOv10-M | 9.03ms |
| YOLOv10-B | 10.04ms |
| YOLOv10-L | 11.52ms |
| YOLOv10-X | 14.67ms |
A.4 Training Cost Analyses
In addition to the inference efficiency analyses, we also investigate the training cost of our YOLOv10 models. We compare with other YOLO variants and measure the training throughput on 8 NVIDIA 3090 GPUs using the official codebases. Tab. 16 presents the comparison results based on the medium model scale. We observe that despite having 500 training epochs, YOLOv10 achieves a high training throughput, making its training cost affordable. We also note that the one-to-many head in the NMS-free training will introduce the extra overhead for YOLOv10. To investigate this, we measure the training cost of YOLOv10 with only the one-to-one head, which is denoted as “YOLOv10-o2o”. As shown in Tab. 16, YOLOv10-M results in a small increase in the training time over “YOLOv10-M-o2o”, about 18s each epoch, which is affordable. To fairly verify the benefit of the one-to-many head in NMS-free training, we also adopt longer 550 training epochs for “YOLOv10-M-o2o”, which leads to a similar training time (29.3 vs. 29.1 hours) but still yields inferior performance (48.9% vs. 51.1% AP) compared with YOLOv10-M.
@@ A.4 Training Cost Analyses
| Model | #Param.(M) | FLOPs(G) | APval(%) | Latency(ms) |
| DEYO-tiny [45] | 4.0 | 8.0 | 37.6 | 2.01 |
| YOLOv10-N | 2.3 | 6.7 | 38.5 | 1.84 |
| DAMO-YOLO-T [71] | 8.5 | 18.1 | 42.0 | 2.21 |
| DAMO-YOLO-S [71] | 16.3 | 37.8 | 46.0 | 3.18 |
| DEYO-S [45] | 14.0 | 26.0 | 45.8 | 3.34 |
| YOLOv10-S | 7.2 | 21.6 | 46.3 | 2.49 |
| DAMO-YOLO-M [71] | 28.2 | 61.8 | 49.2 | 4.57 |
| DAMO-YOLO-L [71] | 42.1 | 97.3 | 50.8 | 6.48 |
| DEYO-M [45] | 33.0 | 78.0 | 50.7 | 7.14 |
| YOLOv10-M | 15.4 | 59.1 | 51.1 | 4.74 |
| YOLOv7 [62] | 36.9 | 104.7 | 51.2 | 17.03 |
| YOLOv10-B | 19.1 | 92.0 | 52.5 | 5.74 |
| YOLOv7-X [62] | 71.3 | 189.9 | 52.9 | 21.45 |
| DEYO-L [45] | 51.0 | 155.0 | 52.7 | 10.00 |
| YOLOv10-L | 24.4 | 120.3 | 53.2 | 7.28 |
| DEYO-X [45] | 78.0 | 242.0 | 53.7 | 15.38 |
| YOLOv10-X | 29.5 | 160.4 | 54.4 | 10.70 |
| Model | #Param.(M) | FLOPs(G) | APval(%) | Latency(ms) |
| YOLOv6-3.0-N [29] | 4.7 | 11.4 | 37.0 | 2.69 |
| Gold-YOLO-N [60] | 5.6 | 12.1 | 39.6 | 2.92 |
| YOLOv10-N (Ours) | 2.3 | 6.7 | 37.7 | 1.84 |
| YOLOv6-3.0-S [29] | 18.5 | 45.3 | 44.3 | 3.42 |
| Gold-YOLO-S [60] | 21.5 | 46.0 | 45.4 | 3.82 |
| YOLO-MS-XS [8] | 4.5 | 17.4 | 43.4 | 8.23 |
| YOLOv10-S (Ours) | 7.2 | 21.6 | 45.6 | 2.49 |
| YOLOv6-3.0-M [29] | 34.9 | 85.8 | 49.1 | 5.63 |
| Gold-YOLO-M [60] | 41.3 | 87.5 | 49.8 | 6.38 |
| YOLOv10-M (Ours) | 15.4 | 59.1 | 50.3 | 4.74 |
| YOLOv6-3.0-L [29] | 59.6 | 150.7 | 51.8 | 9.02 |
| Gold-YOLO-L [60] | 75.1 | 151.7 | 51.8 | 10.65 |
| YOLO-MS [8] | 22.2 | 80.2 | 51.0 | 12.41 |
| YOLOv10-B (Ours) | 19.1 | 92.0 | 51.6 | 5.74 |
| YOLOv10-L (Ours) | 24.4 | 120.3 | 52.4 | 7.28 |
| YOLOv10-X (Ours) | 29.5 | 160.4 | 53.6 | 10.70 |
A.5 More Results on COCO
We measure the latency of YOLOv10 with the original one-to-many training using NMS and report the results on COCO in Tab. 16. Besides, we report the detailed performance of YOLOv10, including AP and AP at different IoU thresholds, as well as AP, AP, and AP across different scales, in Tab. 17. We also present the comparisons with more lightweight detectors, including DAMO-YOLO [71], YOLOv7 [62], and DEYO [45], in Tab. 18. It shows that our YOLOv10 also achieves superior performance and efficiency trade-offs. Additionally, in experiments, we follow previous works [21, 65] to train the models for 500 epochs. We also conduct experiments to train the models for 300 epochs and present the comparison results with YOLOv6 [29], Gold-YOLO [60], and YOLO-MS [8] which adopt 300 epochs, in Tab. 19. We observe that our YOLOv10 also exhibits better performance and inference latency. We also note that despite trained for 500 epochs, YOLOv10 has less training cost compared with these models as presented in Tab. 16.
A.6 Inference Efficiency Comparison on CPU
We present the speed comparison results of YOLOv10 and others on CPU (Intel Xeon Skylake, IBRS) using OpenVINO in Fig. 5. We observe that YOLOv10 also shows state-of-the-art trade-offs in terms of performance and efficiency.
A.7 More Analyses for Holistic Efficiency-Accuracy Driven Model Design
A.8 Visualization Results
@@
A.9 Contribution, Limitation, and Broader Impact